I am confuse how to solve this and make the equation for it please help
1 There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Sawz Logging Co. The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
2 A common mistake when solving equations is the following:
The equation: 2(x _ 2) _ x _ 3
First step in solving: 2x _ 2 _ x _ 3

Answer by maxchuie(2)

300x+225(850-x)=217500
300x + (191250 + -225x) = 217500
191250+300x+-225x=217500
191250+75x=217500
191250 +-191250 +75x = 217500 + -191250
75x=26250
X=350
350 douglas fir
850 -350 = 500 Ponderosa pine